OpenStudy (anonymous):

(1/5 + 9/10)(1/5 - 9/10)

OpenStudy (anonymous):

please...

OpenStudy (anonymous):

help...

OpenStudy (anonymous):

(1/5 + 9/10)(1/5 - 9/10) is a sum and difference problem; (a+b)(a-b)=(a^2-b^2) so (1/5 + 9/10)(1/5 - 9/10)= ((1/25)-(81/100)) (1/25)= (4/100); ((4/100)-(81/100))= -77/100

OpenStudy (anonymous):

Thanks a bunch!

OpenStudy (anonymous):

No problem. Sorry I took so long but when I first typed it OpenStudy made me make an account and Firefox wouldn't allow me to so I switched to Safari.